# Temporal Entanglement Transitions in the Periodically Driven Ising Chain

### v1

@@ Abstract
Periodically driven quantum systems can host non-equilibrium phenomena without static analogs, including in their entanglement dynamics. Here, we discover temporal entanglement transitions in a Floquet spin chain, which correspond to a quantum phase transition in the spectrum of the entanglement Hamiltonian and are signaled by dynamical spontaneous symmetry breaking. We show that these transitions are entanglement-driven, i.e., they require initially entangled states and remain invisible to conventional local observables. Intriguingly, we find these transitions across a broad range of driving frequencies (from adiabatic to high-frequency regime) and independently of drive details, where they manifest as periodic, sharp entanglement spectrum reorganizations marked by the Schmidt-gap closure, a vanishing entanglement echo, and symmetry-quantum-number flips. At high frequencies, the entanglement Hamiltonian acquires an intrinsic timescale decoupled from the drive period, rendering the transitions genuine steady-state features. Finite-size scaling reveals universal critical behavior with correlation-length exponent $\nu=1$, matching equilibrium Ising universality despite its emergence from purely dynamical mechanisms decoupled from static criticality. Our work establishes temporal entanglement transitions as novel features in Floquet quantum matter.
@@
Recent work has begun exploring the interplay between driving and entanglement, including for Floquet-driven conformal field theories [45, 46, 47] and in driven-dissipative systems [48, 49, 50]. Particularly relevant are studies of Page curves in unitary dynamics [51, 52, 53], where entanglement transitions emerge from the competition between entanglement generation and relaxation processes in the presence of conserved charges. While equilibrium quantum phase transitions are intimately connected to entanglement scaling and the structure of the many-body ground state [54, 55, 17], the analogous relationship in driven systems remains unexplored. Since driven systems can access dynamical phases with no equilibrium counterpart, this suggests new entanglement phenomena are waiting to be uncovered. Indeed, developments in geometric Floquet theory [56, 57], advances in conformal field theory approaches to quantum quenches [58, 59, 60], and emerging concepts of entanglement transitions [61, 62, 63, 64, 65, 66, 67] have unearthed novel far-from-equilibrium entanglement dynamics and, in many cases, connections with information-theoretic quantities [68, 69, 70, 71]. However, whether non-analytic entanglement transitions can occur in generic periodically-driven systems remains an open question.
@@
In this Letter, we report the discovery of temporal entanglement transitions via dynamical spontaneous symmetry breaking in the entanglement Hamiltonian (EH) of a periodically driven spin chain. These transitions are uniquely entanglement-driven, manifesting only for initially entangled states and eluding conventional probes such as magnetization or Loschmidt echoes. They manifest as (1) periodic, sharp reorganizations of the entanglement spectrum marked by Schmidt gap closure [72, 73, 74], (2) vanishing overlap between initial and instantaneous entanglement ground states (also known as the entanglement echo [75]), and (3) symmetry quantum number flips [76, 77, 78]. Crucially, these transitions occur across all driving frequencies and exhibit universal critical behavior with correlation length exponent $\nu=1$, matching the equilibrium Ising universality class [79, 80], despite emerging from purely dynamical mechanisms and decoupled from any underlying equilibrium criticality.
@@
Our findings establish temporal entanglement transitions as a distinct class of non-equilibrium phenomenon that is uniquely detectable through entanglement measures while remaining hidden from conventional observables, revealing that the EH itself can undergo quantum phase transitions. At high frequencies, we find that the EH gains an intrinsic timescale decoupled from the drive, rendering its Floquet-periodic transitions as genuine steady-state features. The universal critical behavior persists across different equilibrium phases of the initial state, confirming the transitions’ fundamentally non-equilibrium character, decoupled from the underlying equilibrium phase diagram.
@@
Figure 1: Temporal entanglement transitions in the driven TFIM. (a) Schmidt gap $\Delta\lambda=\lambda_{0}-\lambda_{1}$ closing periodically at critical times. (b) Entanglement echo $|E(t)|^{2}$ vanishing at odd critical times $t_{c}^{(k=\text{odd})}$, as the dominant Schmidt vector $|\lambda_{0}\rangle$ orthogonalizes into a distinct symmetry sector. (c) Subsystem parity expectations showing dynamical spontaneous $\mathbb{Z}_{2}$ symmetry breaking in the entanglement ground state (green) and complementary behavior in the first-excited state (orange). The alternating parity pattern provides consistent physical explanation of the echo’s behavior. Red vertical lines mark critical times of symmetry breaking, and overlaid across all panels to provide a synchronized, consistent picture. Parameters: $L=24$, $L_{A}=9$, $J=1.0$, $h_{0}=2.0$, $\omega=5.0$, $dt=0.01$.
@@
where $J$ is the Ising coupling, $h(t)=(h_{0}/2)\cos(\omega t)$ is the oscillating transverse field, and $\sigma_{i}^{\alpha}$ are Pauli matrices. We initialize the system in the ground state of the static Hamiltonian ${\cal H}_{\text{static}}={\cal H}(t)|_{t=0}$ and evolve under ${\cal H}(t)$ using time-dependent variational principle (TDVP) [81, 82, 83] for matrix product states (MPS) [84], as implemented in the ITensors library [85] (see Supplemental Material (SM) for details [86]). Note that our results are independent of the specific choice of Floquet drive.
@@
A crossing in the two largest Schmidt values, $\lambda_{0}(t)$ and $\lambda_{1}(t)$, signifies a fundamental reorganization of the ES, corresponding to a non-analyticity in the ground state energy $\epsilon_{0}(t)$ of ${\cal H}_{\text{ent}}(t)$ (with fictitious temperature $T_{\text{fict}}=\lim_{n\to\infty}\frac{1}{n}=0$), signaling a temporal quantum phase transition in the EH. Surprisingly, we observe these transitions across all non-zero driving frequencies (from $\omega\ll$ $\min(h_{0},J)$ to $\omega\gg$ $\max(h_{0},J)$), indicating the generic nature of this phenomenon in the periodically-driven TFIM across all frequencies. This behavior emerges specifically when the initial state is entangled (such as the ground state of ${\cal H}_{\text{static}}$), rather than a simple product state or a domain wall state with vanishing entanglement at the subsystem boundary (see SM [86]).
@@
To diagnose dynamical transitions within the entanglement spectrum, we introduce and track specific observables designed to detect symmetry breaking. The central quantity is the entanglement echo $E(t)=\langle\lambda_{0}(0)|\lambda_{0}(t)\rangle$, which measures the fidelity of the instantaneous entanglement ground state to its initial configuration [75]. A vanishing echo signals an orthogonalization of $|\lambda_{0}(t)\rangle$, suggesting a crossing into a distinct symmetry sector. To directly test this, we compute the expectation values of symmetry operators within the entanglement eigenstates. Specifically, we monitor the subsystem parity $\langle\lambda_{n}(t)|P_{A}|\lambda_{n}(t)\rangle$, where $P_{A}=\prod_{i\in A}\sigma_{i}^{x}$. A spontaneous change in the parity of the dominant state $|\lambda_{0}(t)\rangle$, coinciding with a vanishing entanglement echo and a vanishing Schmidt gap $\lambda_{0}-\lambda_{1}$, constitutes the hallmark of a temporal entanglement transition. For each subsequent interval, the dynamics alternate between two dwell times: $T_{-}$, during which the dominant state carries $\langle P_{A}\rangle=-1$, and $T_{+}$, during which it carries $\langle P_{A}\rangle=+1$. At intermediate drives ($\omega\gtrsim$ 7), these intervals become individually regular yet remain unequal, $T_{-}\neq T_{+}$, evidencing partial “Floquet inheritance” by the EH. For higher frequencies ($\omega\gtrsim$ 10), the alternation persists but the intervals synchronize, $T_{-}\approx T_{+}\equiv T_{c}$, and $T_{c}$ saturates to an $\omega$-independent constant value (Fig. 3). This synchronization and saturation indicate complete Floquet inheritance: the EH dynamics are governed by a high-frequency effective (Floquet-Magnus) description, in which the entanglement transitions form a uniformly spaced temporal lattice set by an emergent, drive-induced time-scale rather than the bare period. See Table I and Sec. III.B in SM [86] for further details.
@@
Temporal Entanglement Transitions.— The combined signatures of a temporal entanglement transition are unambiguously observed in the dynamics of the driven TFIM. Fig. 1 presents a comprehensive view of the entanglement dynamics for a subsystem of size $L_{A}=9$, revealing periodic occurrences of the transition.
The onset of a transition is heralded by a critical slowdown of the entanglement spectrum, manifested as a sharp narrowing of the Schmidt gap $\Delta\lambda=\lambda_{0}-\lambda_{1}$ (Fig. 1a). The gap closes to within numerical precision at multiple critical times $t_{c}^{(k)}$}, where we denote the first critical time as $t_{c}^{(k=1)}\equiv t^{*}$, indicating a near-degeneracy of the two largest Schmidt values. This closing gap is the direct signature of an impending energy-level crossing in the EH ${\cal H}_{\text{ent}}(t)$ (recall that the largest and the second largest Schmidt values directly correspond to the ground state and first-excited state energies of ${\cal H}_{\text{ent}}(t)$).
Concurrently, the entanglement echo $E(t)=\left\langle\lambda_{0}(0)\mid\lambda_{0}(t)\right\rangle$ exhibits sharp vanishing at odd critical times (Fig. 1b). At these times $t_{c}^{(k=\text{odd})}$, where the Schmidt gap closes, $|E(t)|^{2}$ drops to zero, establishing that the instantaneous entanglement ground state $|\lambda_{0}(t_{c})\rangle$ orthogonalizes relative to its initial configuration. This confirms the reorganization of quantum correlations and the transition into a distinct symmetry sector of the Hilbert space.
The nature of this sector is revealed by the expectation value of the subsystem parity operator $P_{A}$ (Fig. 1c). For $t<t^{*}$, the ground state of ${\cal H}_{\text{ent}}(t)$ exhibits a well-defined parity of $\langle P_{A}\rangle=+1$. At each transition point $t^{*}$, this parity expectation flips discontinuously to $\langle P_{A}\rangle=-1$, indicating a dynamical spontaneous breaking of the $\mathbb{Z}_{2}$ symmetry by the entanglement ground state. Notably, the first excited state $|\lambda_{1}(t)\rangle$ displays the opposite parity, confirming that the gap closing corresponds to an exchange of roles between two symmetry-broken states. The subsystem magnetization $\langle M_{A}\rangle$ [80] and Loschmidt echo rate function [87, 88, 89, 90] (verified in SM [86]) remain smooth through these events, confirming the transition is of a purely entanglement-driven, dynamical spontaneous symmetry-breaking character.
@@
We observe these transition signatures across all non-zero driving frequencies $\omega$ studied, indicating the generic nature of this phenomenon in periodically driven Ising chain. This behavior emerges specifically when the initial state is entangled (such as the ground state of ${\cal H}_{\text{static}}$), rather than a simple non-entangled state (see SM [86]). While transitions occur at all frequencies, the periodic regularity of critical times $t_{c}^{(k)}$} becomes pronounced only at higher driving frequencies, suggesting a crossover to Floquet-like behavior at the level of the EH (details provided in the SM [86]).
@@
The synchronization of these three diagnostics, namely the closing of the Schmidt gap, the vanishing of the entanglement echo, and the discontinuous flip in parity, provides definitive evidence of a temporal entanglement transition. These are not independent events but are intrinsically linked manifestations of the same underlying phenomenon: a periodically-driven, dynamical quantum phase transition of the EH ${\cal H}_{\text{ent}}(t)$.
@@
Finite-Size Scaling.— The temporal entanglement transitions are marked by non-analytic kinks in the ground state energy $\epsilon_{0}(t)=-\ln\lambda_{0}(t)$ of the EH. To establish the critical nature and universality of temporal entanglement transitions, we perform comprehensive finite-size scaling analysis across subsystem sizes $L_{A}=4,5,\ldots,12$.
@@
Fig. 2 demonstrates the scaling properties of the first critical time $t_{c}^{(1)}=t^{*}$ and the minimum Schmidt gap density $S_{\text{min}}/L_{A}$ at criticality. The critical time exhibits power-law scaling $t^{*}/L_{A}\propto L_{A}^{-1/\nu}$ with critical exponent $\nu\simeq 1.00$ (Fig. 2a) that corresponds to the divergence of correlation length. Simultaneously, the critical entropy density follows $S_{\text{min}}/L_{A}\propto L_{A}^{-a}$ with $a\simeq 1$ (Fig. 2b). Surprisingly, the correlation length exponent $\nu=1$ establishes this as a continuous quantum phase transition for all ranges of driving frequencies, belonging to the same universality class as the equilibrium 2D classical Ising/1D quantum TFIM. Details provided in the SM [86] confirm that this exponent match is coincidental: to the best of our knowledge, temporal entanglement transitions appear to occur independently of the equilibrium phase diagram.
@@
The universality of these transitions is demonstrated through data collapse using the scaling ansatz (recall $\epsilon_{0}=S_{\text{min}}$):
$\frac{\epsilon_{0}}{L_{A}}=\frac{1}{L_{A}^{a}}\mathcal{F}\left[\left(\frac{t}{L_{A}}-\frac{t_{c}}{L_{A}}\right)L_{A}^{1/\nu}\right],$ (2)
@@
where $\mathcal{F}$ is a universal scaling function, and $\nu$ is the critical exponent corresponding to diverging correlation length. Fig. 2c shows excellent collapse of data from all subsystem sizes onto a universal curve, validating the scaling hypothesis. The agreement is further verified in Fig. 2d, which shows the raw data near the critical point.
@@
Significantly, the same exponent describes all subsequent kinks at $t_{c}^{(k\geq 1)}$ for higher driving frequencies ($\omega\gtrsim 5.0$ for our parameter choices), where the EH inherits Floquet-like periodicity. This universality across multiple transitions suggests a common underlying fixed point. The exponent $\nu\simeq 1.00$ defines a new universality class for non-equilibrium entanglement dynamics in driven systems, which appears to be completely decoupled from the underlying equilibrium criticality (more on this below).
@@
Figure 2: Finite-size scaling analysis of temporal entanglement transitions. (a) Critical time scaling $t^{*}/L_{A}\propto L_{A}^{-1/\nu}$ with $\nu=1.00$. (b) Critical entropy density scaling $\epsilon_{0}/L_{A}\propto L_{A}^{-a}$ with $a=1.00$. (c) Universal data collapse using scaling ansatz in Eq. (2). (d) Raw data before scaling collapse. The exponent $\nu=1$ establishes the same universality class for our driven non-equilibrium entanglement spectrum as equilibrium 2D classical Ising/1D TFIM. Parameters: $L=24$, $J=1.0$, $h_{0}=2.0$, $\omega=5.0$, $dt=0.01$.
@@
Figure 3: Frequency dependence of the first critical time $t^{*}$ across subsystem sizes $L_{A}=4-12$. At low frequencies, $t^{*}\propto\omega^{-1}$ (adiabatic regime), while at high frequencies ($\omega\gtrsim 10$), $t^{*}$ saturates to frequency-independent values determined solely by subsystem size, indicating a crossover to Floquet steady-state behavior where the EH develops an intrinsic timescale. The universal saturation indicates that temporal entanglement transitions become intrinsic properties of the effective time-independent dynamics rather than driven phenomena, validating the Floquet-Magnus description and establishing characteristic timescales independent of the external driving protocol. Parameters: $L=24$, $J=1.0$, $h_{0}=2.0$. Time steps are chosen as $dt=0.01$ for $\omega=[0.1,10.0]$, $dt=0.002$ for $\omega=\{30,50\}$ and $dt=0.001$ for $\omega=\{70,100\}$. A total of 153 data points in this plot show excellent collapse (see Fig. 9 of SM [86]).
@@
Effective Steady State at High Frequency.— At high driving frequencies, the temporal entanglement transitions can be understood through Floquet-Magnus theory [91, 92, 93]. We decompose the time-dependent Hamiltonian as $\mathcal{H}(t)=\mathcal{A}+\mathcal{B}\cos(\omega t)$, where $\mathcal{A}=-J\sum_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}$ and $\mathcal{B}=-(h_{0}/2)\sum_{i}\sigma_{i}^{x}$. The effective time-independent Hamiltonian emerges as a systematic expansion $\mathcal{H}_{\text{eff}}=\mathcal{H}_{0}+\mathcal{H}_{1}+\mathcal{H}_{2}+\cdots$ ${\cal H}_{(n)}$ is of order ${\cal O}(\omega^{-n})$], where the leading corrections are

### v2

@@ Abstract
Periodically driven quantum systems can host non-equilibrium phenomena without static analogs, including in their entanglement dynamics. Here, we discover temporal entanglement transitions (TET) in a Floquet spin chain, which correspond to a quantum phase transition in the spectrum of the entanglement Hamiltonian and are signaled by dynamical spontaneous symmetry breaking. We identify the symmetry principles underlying these transitions: they appear when the driven Hamiltonian preserves global symmetry (here, $\mathbb{Z}_{2}$), the initial state respects this symmetry, and the reduced density matrix carries weight in both subsystem-parity sectors, with TET occurring precisely when the sector weights become equal (given the previous two conditions are also satisfied). Intriguingly, we find these transitions across a broad range of driving frequencies (from adiabatic to high-frequency regime) and independently of drive details, where they manifest as periodic, sharp entanglement spectrum reorganizations marked by the Schmidt-gap closure, a vanishing entanglement echo, and symmetry-quantum-number flips, while remaining invisible to conventional local observables. At high frequencies, the entanglement Hamiltonian acquires an intrinsic timescale decoupled from the drive period, rendering the transitions genuine steady-state features. Finite-size scaling reveals universal critical behavior with correlation-length exponent $\nu=1$, matching equilibrium Ising universality despite its emergence from purely dynamical mechanisms decoupled from static criticality. Our work establishes TET as novel features in Floquet quantum matter.
@@
Recent work has begun exploring the interplay between driving and entanglement, including for Floquet-driven conformal field theories [45, 46, 47, 48] and in driven-dissipative systems [49, 50, 51]. Particularly relevant are studies of Page curves [52, 53, 54], where entanglement transitions emerge from competition between entanglement generation and relaxation in the presence of conserved charges. While equilibrium quantum phase transitions are intimately connected to entanglement scaling and the structure of the many-body ground state [55, 56, 17], the analogous relationship in driven systems remains unexplored. Since driven systems can access dynamical phases with no equilibrium counterpart, this suggests new entanglement phenomena waiting to be uncovered. However, whether non-analytic entanglement transitions can occur in generic periodically-driven systems remains an open question.
@@
In this Letter, we report the discovery of temporal entanglement transitions (TET) via dynamical spontaneous symmetry breaking in the entanglement Hamiltonian (EH) of a periodically driven spin chain. These transitions require $\mathbb{Z}_{2}$-symmetric driven Hamiltonian and initial state, with critical times occurring at sector-weight degeneracies (made precise below), and manifest through three synchronized signatures: Schmidt gap closure [57, 58, 59], vanishing entanglement echo [60], and parity flips [61, 62, 63], while remaining invisible to conventional observables (magnetization, Loschmidt echo). Crucially, these transitions occur across all driving frequencies and exhibit universal critical behavior with correlation length exponent $\nu=1$, matching the equilibrium Ising universality class [64, 65], despite emerging purely dynamically, decoupled from any underlying equilibrium criticality.
@@
Our findings establish TET as a distinct class of non-equilibrium phenomenon that is uniquely detectable through entanglement measures while remaining hidden from conventional observables, revealing that the EH itself can undergo quantum phase transitions. At high frequencies, we find that the EH gains an intrinsic timescale decoupled from the drive, rendering its Floquet-periodic transitions as genuine steady-state features. The universal critical behavior persists whenever the above Z2-symmetry condition and sector-weight equality are satisfied (including across different equilibrium phases that preserve Z2), confirming a fundamentally non-equilibrium character decoupled from the underlying equilibrium phase diagram (see Supplemental Material (SM) [66] for details).
@@
Figure 1: Temporal entanglement transitions (TET) in the driven TFIM. (a) Schmidt gap $\Delta\lambda=\lambda_{0}-\lambda_{1}$ closing periodically at critical times. (b) Entanglement echo $|E(t)|^{2}$ vanishing at odd critical times $t_{c}^{(k=\text{odd})}$, as the dominant Schmidt vector $|\lambda_{0}\rangle$ orthogonalizes into a distinct symmetry sector. (c) Subsystem parity expectations showing dynamical spontaneous $\mathbb{Z}_{2}$ symmetry breaking in the entanglement ground state (green) and complementary behavior in the first-excited state (orange). The alternating parity pattern physically underlies the echo’s behavior. Red vertical lines mark critical times of symmetry breaking, and are overlaid across all panels to provide a synchronized, consistent picture. Parameters: $L=24$, $L_{A}=9$, $J=1.0$, $h_{0}=2.0$, $\omega=5.0$, $dt=0.01$.
@@
where $J$ is the Ising coupling, $h(t)=(h_{0}/2)\cos(\omega t)$ is the oscillating transverse field, and $\sigma_{i}^{\alpha}$ are Pauli matrices. We initialize the system in the ground state of the static Hamiltonian ${\cal H}_{\text{static}}={\cal H}(t)|_{t=0}$ and evolve under ${\cal H}(t)$ using time-dependent variational principle (TDVP) [67, 68, 69] for matrix product states (MPS) [70], as implemented in the ITensors library [71] (see SM [66]). Note that our results are independent of the specific Floquet drive.
@@
TET are a universal consequence at all driving frequencies provided (i) the driven Hamiltonian and (ii) the initial state preserve the global $\mathbb{Z}_{2}$ symmetry, and (iii) $\rho_{A}(t)$ has support in both subsystem-parity sectors. We quantify (iii) by the sector weights $w_{\pm}(t)=\operatorname{Tr}[P_{\pm}\rho_{A}(t)]$, with $P_{\pm}=(\mathds{1}_{A}\pm P_{A})/2$ and $P_{A}=\prod_{i\in A}\sigma_{i}^{x}$. Conditions (i)–(ii) imply $\rho_{A}(t),P_{A}]=0$, which we verify using $C(t)=\|[\rho_{A}(t),P_{A}]\|$. Within this symmetry-preserving regime, the critical times $t_{c}^{(k)}$ are selected by the operational criterion $w_{+}(t_{c}^{(k)})=w_{-}(t_{c}^{(k)})$ (within the evolution window), at which the Schmidt gap closes and the leading Schmidt vectors exchange subsystem-parity discontinuously. By contrast, an isolated suppression of the Schmidt gap or a smooth parity drift—even if accompanied by a sector-weight crossing—indicates a rounded crossover rather than a TET (see Sec. III of the SM [66]).
@@
A crossing in the two largest Schmidt values, $\lambda_{0}(t)$ and $\lambda_{1}(t)$, signifies a fundamental reorganization of the ES, corresponding to a non-analyticity in the ground state energy $\epsilon_{0}(t)$ of ${\cal H}_{\text{ent}}(t)$ (with fictitious temperature $T_{\text{fict}}=\lim_{n\to\infty}\frac{1}{n}=0$), signaling a sharp temporal quantum phase transition in the EH. Surprisingly, we observe this behavior from the adiabatic to the high-frequency regime. This behavior emerges when the global $\mathbb{Z}_{2}$ symmetry is preserved by the driven Hamiltonian and respected by the initial state, with critical times occurring at sector-weight crossings $w_{+}(t_{c}^{(k)})=w_{-}(t_{c}^{(k)})$. Conversely, $\mathbb{Z}_{2}$-breaking initial states (including symmetry-broken ferromagnetic ground states as well as product states such as domain wall and random product states) do not exhibit TET, independent of the drive protocol (see Sec. III in SM [66]).
@@
To diagnose dynamical transitions within the entanglement spectrum, we introduce and track specific observables designed to detect symmetry breaking. The central quantity is the entanglement echo $E(t)=\langle\lambda_{0}(0)|\lambda_{0}(t)\rangle$, which measures the fidelity of the instantaneous entanglement ground state to its initial configuration [60]. A vanishing echo signals an orthogonalization of $|\lambda_{0}(t)\rangle$, suggesting a crossing into a distinct symmetry sector. To directly test this, we compute the expectation values of symmetry operators within the entanglement eigenstates. Specifically, we monitor the subsystem parity $\langle\lambda_{n}(t)|P_{A}|\lambda_{n}(t)\rangle$. A spontaneous change in the parity of the dominant state $|\lambda_{0}(t)\rangle$, coinciding with a vanishing entanglement echo and a vanishing Schmidt gap $\lambda_{0}-\lambda_{1}$, constitutes the hallmark of a TET. For each subsequent interval, the dynamics alternate between two dwell times: $T_{-}$, during which the dominant state carries $\langle P_{A}\rangle=-1$, and $T_{+}$, during which it carries $\langle P_{A}\rangle=+1$. At intermediate drives ($\omega\gtrsim$ 7), these intervals become individually regular yet remain unequal, $T_{-}\neq T_{+}$, evidencing partial “Floquet inheritance” by the EH. For higher frequencies ($\omega\gtrsim$ 10), the alternation persists but the intervals synchronize, $T_{-}\approx T_{+}\equiv T_{c}$, and $T_{c}$ saturates to an $\omega$-independent constant value (Fig. 3). This synchronization and saturation indicate complete Floquet inheritance: the EH dynamics are governed by a high-frequency effective (Floquet-Magnus) description, in which the entanglement transitions form a uniformly spaced temporal lattice set by an emergent, drive-induced time-scale rather than the bare period. See Table II, Sec. III and Sec. IV.B in SM [66] for further details.
@@
Temporal Entanglement Transitions.— The combined signatures of a TET are unambiguously observed in the driven TFIM (Fig. 1, $L_{A}=9$). The Schmidt gap $\Delta\lambda=\lambda_{0}-\lambda_{1}$ (Fig. 1a) closes to numerical precision at critical times $t_{c}^{(k)}$ (with first critical time $t_{c}^{(k=1)}\equiv t^{*}$), signaling energy-level crossing in ${\cal H}_{\text{ent}}(t)$. Concurrently, the entanglement echo $E(t)=\langle\lambda_{0}(0)\mid\lambda_{0}(t)\rangle$ (Fig. 1b) vanishes at odd $t_{c}^{(k=\text{odd})}$, confirming that $|\lambda_{0}(t_{c})\rangle$ orthogonalizes into a distinct symmetry sector. The subsystem parity $\langle P_{A}\rangle$ (Fig. 1c) flips discontinuously at each $t_{c}^{(k)}$ between +1+1 and 1-1, evidencing dynamical spontaneous $\mathbb{Z}_{2}$ breaking with respect to the largest Schmidt vector $|\lambda_{0}(t)\rangle$, while the first excited state $|\lambda_{1}(t)\rangle$ carries opposite parity.
@@
We observe these signatures across all non-zero driving frequencies, indicating the generic nature of this phenomenon in the periodically driven Ising chain. While the transitions remain sharp, the spacing of critical times $t_{c}^{(k)}$} becomes nearly periodic only at higher driving frequencies, indicating a crossover in the timing (Floquet inheritance) rather than in the transition itself (see SM [66]).
@@
The synchronization of these three diagnostics: the closing of the Schmidt gap, the vanishing of the entanglement echo, and the discontinuous flip in parity, provides definitive evidence of a TET. These are not independent events but are intrinsically linked manifestations of the same underlying phenomenon: a periodically-driven, dynamical quantum phase transition of the EH ${\cal H}_{\text{ent}}(t)$.
@@
Finite-Size Scaling.— The TET are marked by non-analytic kinks in the ground state energy $\epsilon_{0}(t)=-\ln\lambda_{0}(t)$ of the EH. To establish the critical nature and universality of TET for the critical initial state (where all three conditions are robustly satisfied in the short-time dynamics), we perform comprehensive finite-size scaling analysis across subsystem sizes $L_{A}=4-12$.
@@
Fig. 2 demonstrates the scaling properties of the first critical time $t_{c}^{(1)}=t^{*}$ and the minimum Schmidt gap density $S_{\text{min}}/L_{A}$ at criticality. The critical time exhibits power-law scaling $t^{*}/L_{A}\propto L_{A}^{-1/\nu}$ with critical exponent $\nu\simeq 1.00$ (Fig. 2a) that corresponds to the divergence of correlation length. Simultaneously, the critical entropy density follows $S_{\text{min}}/L_{A}\propto L_{A}^{-a}$ with $a\simeq 1$ (Fig. 2b). Surprisingly, the correlation length exponent $\nu=1$ establishes this as a continuous quantum phase transition for all ranges of driving frequencies, belonging to the same universality class as the equilibrium 2D classical Ising/1D quantum TFIM. Note that this exponent match $\nu=1$ appears coincidental; TET seem to emerge independently of the equilibrium phase diagram, as confirmed in SM [66].
@@
The universality of these transitions for the critical initial state is demonstrated through data collapse using the scaling ansatz (recall $\epsilon_{0}=S_{\text{min}}$):
$\frac{\epsilon_{0}}{L_{A}}=\frac{1}{L_{A}^{a}}\mathcal{F}\left[\left(\frac{t}{L_{A}}-\frac{t^{*}}{L_{A}}\right)L_{A}^{1/\nu}\right],$ (2)
@@
where $\mathcal{F}$ is a universal scaling function, and $\nu$ is the critical exponent corresponding to diverging correlation length. For our critical initial state, $t^{*}$ occurs within the simulated time window for all driving frequencies studied, enabling a uniform finite-size scaling analysis from the adiabatic to the high-frequency regime. Fig. 2c shows excellent collapse of data from all subsystem sizes onto a universal curve, validating the scaling hypothesis. The agreement is further verified in Fig. 2d, which shows the raw data near the critical point.
@@
Significantly, the same exponent describes all subsequent kinks at $t_{c}^{(k\geq 1)}$ for higher driving frequencies ($\omega\gtrsim 5.0$ for our parameter choices), where the EH inherits Floquet-like periodicity. This universality across multiple transitions suggests a common underlying fixed point. For the critical initial state, the exponent $\nu\simeq 1.0$ defines a new universality class for non-equilibrium entanglement dynamics in driven systems, which appears to be completely decoupled from the underlying equilibrium criticality (more on this below).
@@
Figure 2: Finite-size scaling analysis of TET starting with critical initial state $(h_{0}/2=1.0)$. (a) Critical time scaling $t^{*}/L_{A}\propto L_{A}^{-1/\nu}$ with $\nu=1.00$. (b) Critical entropy density scaling $\epsilon_{0}/L_{A}\propto L_{A}^{-a}$ with $a=1.00$. (c) Universal data collapse using scaling ansatz in Eq. (2). (d) Raw data before scaling collapse. The exponent $\nu=1$ establishes the same universality class for our driven non-equilibrium entanglement spectrum as equilibrium 2D classical Ising/1D TFIM. Parameters: $L=24$, $J=1.0$, $h_{0}=2.0$, $\omega=5.0$, $dt=0.01$.
@@
Figure 3: Frequency dependence of the first critical time $t^{*}$ across subsystem sizes $L_{A}=4-12$ for critical initial state $(h_{0}/2=1.0)$. At low frequencies, $t^{*}\propto\omega^{-1}$ (adiabatic regime), while at high frequencies ($\omega\gtrsim 10$), $t^{*}$ saturates to frequency-independent values determined solely by subsystem size, indicating a crossover to Floquet steady-state behavior where the EH develops an intrinsic timescale. The universal saturation indicates that TET become intrinsic properties of the effective time-independent dynamics rather than driven phenomena, validating the Floquet-Magnus description and establishing characteristic timescales independent of the external driving protocol. Parameters: $L=24$, $J=1.0$, $h_{0}=2.0$. Time steps are chosen as $dt=0.01$ for $\omega=[0.1,10.0]$, $dt=0.002$ for $\omega=\{30,50\}$ and $dt=0.001$ for $\omega=\{70,100\}$. A total of 153 data points in this plot show excellent collapse (see Fig. 9 of SM [66]).
@@
Effective Steady State at High Frequency.— At high driving frequencies, TET acquire an intrinsic timescale explained by Floquet-Magnus theory [72, 73, 74]. We decompose the time-dependent Hamiltonian as $\mathcal{H}(t)=\mathcal{A}+\mathcal{B}\cos(\omega t)$, where $\mathcal{A}=-J\sum_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}$ and $\mathcal{B}=-(h_{0}/2)\sum_{i}\sigma_{i}^{x}$. The effective time-independent Hamiltonian emerges as a systematic expansion $\mathcal{H}_{\text{eff}}=\mathcal{H}_{0}+\mathcal{H}_{1}+\mathcal{H}_{2}+\cdots$ ${\cal H}_{(n)}$ is of order ${\cal O}(\omega^{-n})$], where the leading corrections are